OpenStudy (anonymous):

The trigonometric form of -5+4i is ?

myininaya (myininaya):

you mean polar form?

OpenStudy (anonymous):

r(cos theta+ i sin theta)

myininaya (myininaya):

\[x+yi=-5+4i\] => x=-5 y=4 \[r=\sqrt{x^2+y^2}\] We do \[\theta=\tan^{-1}(\frac{y}{x})+180\] Since \[\theta\] is in the second quadrant

OpenStudy (anonymous):

so r=√41 i think

OpenStudy (anonymous):

why do you add 180 ?

OpenStudy (anonymous):

so i got theta = -38

myininaya (myininaya):